Reports 1-1 of 1 Clear search Modify search
AdV-SAT (Suspension control upgrade)
boschi, baratti, dandrea, palaia, piendibene, razzano, vacatello - 18:40 Tuesday 18 March 2025 (66399) Print this report
Installation of new PDU switch

Today we installed at the NE a new PDU switch between the smart sockets and the stepping motor units to reduce the 50Hz electrical noise from the motors themselves. The switch has been physically placed below the smart socket.

While doing so, the Sc crate unintendedly switched off, so we took the chance to also replace its malfunctioning cooling fans.

After the intervention, we closed all the control loops without any major issue. We confirmed that all motors are properly working with the same interface as before, and that turning off a motor from the interface also switches off the corresponding PDU switch.

Comments to this report:
bersanetti - 15:46 Friday 28 March 2025 (66403) Print this report

This entry is dated 18/03 but it got swallowed back in the Drafts.

The intervention from this morning, done with the North Arm still locked when the crate switched off (Fig. 1), had three consequences:

  • excitation of the NE violin modes (Fig. 2, LOW_NOISE_2 today and two weeks ago);
  • excitation of the NI vertical mode at 6.9 Hz (Fig. 3, CARM_NULL_1F today and last night), damped by hand in the afternoon as it saturated the B1 dynamics in DC Readout and above;
  • an increase of the Tolm errors of the NI DSP (Fig. 4), maybe due to a jump on the NE flag (Fig. 5)?
Images attached to this comment
Search Help
×

Warning

×